# SODAs: Sparse Optimization for the Discovery of Differential and Algebraic Equations

### v1

@@ 2.3 Benefits and challenges of model discovery in the DAE framework
In addition to the high data quality requirements, a fundamental issue is that implicit-SINDy and PI-SINDy methods do not identify the algebraic variables and associated equations within a DAE system. The algebraic constraints in the DAEs typically carry physical significance related to physical constraints or quasi-steady-state approximations, and this intuitive meaning is lost in the equivalent ODE formulation. Creating an explicit system of ODEs in the form (8) can be complex, computationally intensive, or even impossible for higher-index DAE systems or when the algebraic constraints are not differentiable in time [39]. Furthermore, small parameter changes in the original DAE system often lead to distinct reduced ODE forms. Solving or discovering DAEs directly preserves the underlying structure of the problem. An additional benefit of identifying the algebraic relationships is that the library can be iteratively simplified by removing an algebraic variable for every algebraic equation discovered. Removal of terms with algebraic variables reduce multicollinearity among the candidate library terms due to the explicit algebraic constraints, improving the conditioning of the library. Consequently, the discovery of DAEs provides a more interpretable and computationally beneficial framing.
Discovering a DAE system in its semi-explicit form (2)–(3) also offers flexibility for numerical analysis and simulation. For example, in CRN applications, simulations are often carried out using ODE solvers designed for stiff systems; however, in applications involving high-index DAEs, solving a reduced ODE system can lead to numerical instability [39, 40, 34]. In these cases, single-step and multi-step methods for solving DAEs in their non-reduced forms are preferred [39, 40, 34, 41, 35]. By first discovering the DAE system in the semi-explicit form, the user can choose either to convert it to a reduced ODE or to keep the original semi-explicit form, depending on which numerical solvers are most appropriate. In comparison, converting a reduced ODE system back into a DAE system with explicit constraints is often not feasible.
To address all the issues discussed so far—including the interpretability of physical constraints, the numerical instabilities caused by perfect multicollinearity in the candidate library, and the adversial effects of noise present in implicit ODE discovery—we propose the method SODAs. By separating the discovery of differential and algebraic equations, we directly identify the algebraic variables and equations and eliminate the impact of noisy derivatives on the discovery of algebraic equations. We can then utilize the knowledge gained from the discovered algebraic equations to better condition the library before attempting to discover differential equations.
@@ 3 Methods: Sparse Optimization for Differential-Algebraic Systems
Figure 1: SODAs algorithm explained using a chemical reaction network example. (a) Collect time series and construct candidate library matrix. (b) Algebraic Relation Finder: iteratively finds algebraic relationships and refines the candidate library based on the discovered relationship. In each iteration $k$, the best relationship is selected based on its score $S^{k}$, and a complexity score is assigned to each term within this relationship. This complexity score is used to select a term to refine the candidate library for the next iteration. (c) Dynamic Finder: the refined candidate library from the algebraic finder step is used to find the system of ODEs. (d) Assembled DAEs: The algebraic and differential equations from the algebraic and dynamic finder steps respectively, are assembled to form the final set of DAEs (left), which is equivalent to the true set of DAEs (right).
In this section, we introduce our algorithm SODAs (Figure 1, Algorithm 1), and demonstrate how it can be used to discover DAEs in their explicit form (2)–(3). The core idea is to separate the discovery of the algebraic and dynamic state equations. DaeFinder is a Python implementation of SODAs that allows users to easily create popular candidate libraries relevant to many applications and efficiently implement various components of the SODAs algorithm. Through DaeFinder, SODAs iteratively identifies algebraic relationships, removes them from the library, and reduces multicollinearity within the library. After this process, SINDy is applied to discover the dynamic equations from the remaining library terms. We illustrate the application of SODAs using an example of a simple CRN (4)–(7) (Figure 1a).
@@
Let $\mathbf{X}=[\mathbf{X}_{1}\;\mathbf{X}_{2}\;\ldots\;\mathbf{X}_{d}]$ denote the $N\times d$ time-series data matrix, where each column $\mathbf{X}_{i}$ contains possibly noisy measurements of the state $x_{i}$ at time points $\{t_{1}=0,\ldots,t_{N}=T\}$, such that $\mathbf{X}_{ij}\approx x_{j}(t_{i})$. Note that the measurements $x_{i}(t)$ may also obtained through transformations of the original state variables. However, for simplicity, throughout this paper, we assume that all states are directly observable with some measurement noise. SODAs is typically applied to DAE systems where the terms appearing in equations (2)–(3) can be written as linear combination of a set of predefined library terms which comes from prior knowledge. We start by defining a candidate library $\Theta=\{\theta_{1}(\mathbf{x}),\ldots,\theta_{J}(\mathbf{x})\}$, which is a set of functions of the state variables. Based on the time-series data $\mathbf{X}$, we next form the initial candidate library matrix denoted by $\Theta_{0}=[\theta_{1}(\mathbf{X})\ \theta_{2}(\mathbf{X})\ \ldots\ \theta_{J}(\mathbf{X})]$, which is an $N\times J$ matrix where the $j^{\text{th}}$ column represents the evaluation of the candidate function $\theta_{j}$ at time points $\{t_{1}=0,\ldots,t_{N}=T\}$. The DaeFinder package provides various choices for library functions and their corresponding matrix creation, such as polynomial monomial libraries, Fourier series, phase-difference terms, and various custom functions that can be controlled using a connectivity matrix (which is useful in applications like power-grid networks where pre-existing knowledge about connections between nodes can guide library construction). We have also developed the package to interface with the PySINDy package [42] such that PySINDy’s library construction features can also be used with DaeFinder. In our applications, a polynomial library is used for the CRN and the nonlinear pendulum, while a combination of polynomial and trigonometric functions is used for the power grid. After forming the candidate library, SODAs consists of two major steps: the algebraic finder step (Figure 1b) and the dynamical system finder step (Figure 1c).
@@
For many systems, it is unclear a priori which states are differential variables and which are algebraic, especially if separation of timescales is expected and a subset of variables are in quasi-steady state. Our approach does not necessarily require pre-existing knowledge about which states are in the algebraic relations. Notwithstanding, prior knowledge on state variables that take part in algebraic equations can be incorporated into the pipeline. For unknown algebraic equations, the algebraic finder step involves iteratively identifying algebraic relationships and reducing multicollinearity in the library by removing these relationships. Once all algebraic equations are found, the remaining dynamical equations are identified from the refined library using an existing method like SINDy [10, 18, 16] or IDENT [29, 30].
@@
$L_{l}(\mathbf{p}^{l})$ $=\sum_{r=1}^{N}\left\|\theta_{l}(\mathbf{x}(t_{r}))-\sum_{\begin{subarray}{c}j\in\mathcal{I}_{k}\\ j\neq l\end{subarray}}p_{j}^{l}\theta_{j}(\mathbf{x}(t_{r}))\right\|^{2}+\alpha\|\mathbf{p}^{l}\|_{0},\quad l\in\mathcal{I}_{k-1}.$ (13)
@@
where $\Xi=\left[\mathbf{p}^{1}\ \mathbf{p}^{2}\ \ldots\ \mathbf{p}^{J_{k}}\right]$ with constraints $\Xi_{jj}=0$ and $\|\Xi\|_{0}$ is minimized (Figure 1b, left). Note that, in practice, the matrix form is not used, but instead each $\mathbf{p}^{l}$ is optimized in parallel by minimizing the loss function $L_{l}(\mathbf{p}^{l})$. A similar parallelized approach is introduced in [27] for semi-combinatorial discovery of implicit equations of the form (11), whereas in SODAs this parallel fitting approach is used for finding algebraic relationships. Notably, derivatives are not involved in the algebraic finder step and do not need to be estimated from the noisy data. There are multiple established approaches to enforce the sparsity of the parameter matrix $\Xi$. The DaeFinder package supports popular alternatives to $\ell_{0}$ regularization, including LASSO and various sequential-thresholding methods combined with regularizers like Ridge regression and Elastic Net. Previous work has shown that soft thresholded ordinary least squares estimator (STOLS) outperforms the usage of LASSO alone for SINDy [43]. Accordingly, we employed LASSO followed by sequential thresholding as the default solver in our numerical experiments, unless specified otherwise.
@@
The next step of the algorithm is to select the best-fitting algebraic relationship across all the candidate fits in (13) (Figure 1b, right). To identify the most likely relationship, we assign a score $S^{k}_{l}$ to each candidate fit in (13). In our examples, we have used the coefficient of determination $\mathcal{R}^{2}$ [44], which ranges from 0 to 1, where larger values indicate a good fitting. Other criteria, such as AIC and BIC (see [45, 46]), that balance training error with model simplicity can also be used to define $S^{k}_{l}$. The fit with the highest score is then identified as the most prominent algebraic relationship in the library $\Theta_{k-1}$ and is further used to refine $\Theta_{k-1}$. In our experience, conservation laws that involve first-order terms are initially identified as the best relationships during this process.
@@
After identifying the best algebraic relationship based on the score $S^{k}_{l}$ for the $k^{\text{th}}$ iteration, any common factors in that relationship are factored out to obtain a reduced equation $g_{k}(\mathbf{x})=0$. To address the multicollinearity introduced by this algebraic relationship, we remove one of the terms in $g_{k}(\mathbf{x})$. To decide which term to remove, each term $\theta^{(k)}$ in $g_{k}(\mathbf{x})$ is assigned a complexity score (Figure 1b, right). In polynomial libraries, this score is set to be the degree of the term by default, i.e., higher-degree monomials should have a higher complexity score. The term with the highest complexity score, $\theta^{(k)}_{\text{high }}$ (circled in Figure 1b, right), is then removed from the library. The rationale behind this refinement is that if $\theta^{(k)}_{\text{high }}$ is part of the discovered algebraic relationship $g_{k}(\mathbf{x})=0$, then $\theta^{(k)}_{\text{high }}$ can be expressed as a linear combination of the other terms in $g_{k}(\mathbf{x})$, so there is no loss of information by removing a term from the library. By keeping lower rather than higher complexity terms in the library, we reduce noise amplification [47]. If multiple terms share the same complexity score, then either expert knowledge or quality of measurement of states (through statistical measures like variance) can also be used to choose one term over the other, or a term can be selected randomly. Additionally, for any $\theta\in\Theta_{k-1}$, the equation $\theta\,g_{k}(\mathbf{x})=0$ also represents a potential algebraic relationship that can lead to perfect multicollinearity. To remove all degenerate relationships of the form $\theta\,g_{k}(\mathbf{x})=0$, all multiples of $\theta^{(k)}_{\text{high }}$are also removed from $\Theta_{k-1}$. Let $\Gamma_{k}$ be a subset of the candidate library matrix $\Theta_{k-1}$ that has $\theta^{(k)}_{\text{high }}$ as a factor; then $\Theta_{k}=\Theta_{k-1}\setminus\Gamma_{k}$, where all elements of $\Gamma_{k}$ are removed from $\Theta_{k-1}$ as part of the refinement process. It follows that the libraries at each refinement step form a decreasing chain, $\Theta_{k}\subset\Theta_{k-1}\subset\cdots\subset\Theta_{0}.$
Mitigating perfect multicollinearity in the candidate library
@@
As an example, in the CRNs (4)–(7), the algebraic relationship $[E]+[ES]-E_{0}=0$ can cause the following perfect multicollinearity in the candidate library matrix $\Theta$ that has up to degree 2 monomials:
$[E]+[ES]-E_{0}=0,$ (15)
$[E]\left([E]+[ES]-E_{0}\right)=0,$ (16)
$[S]\left([E]+[ES]-E_{0}\right)=0,$ (17)
$[P]\left([E]+[ES]-E_{0}\right)=0,$ (18)
$[ES]\left([E]+[ES]-E_{0}\right)=0.$ (19)
@@
For higher-order candidate libraries, there will be more relationships. All such degenerate algebraic relationships make up the nullspace of the library matrix $\Theta$. If any one of the degenerate equations (16)–(19) is identified as the relationship with the best $S_{l}^{k}$, it is first reduced to the form (15). Both terms in this reduced relationship (i.e., $[E]$ and $[ES]$) have the same complexity score of 1 based on the degree of the monomial (Figure 1b, top-right). In the absence of any other information, we randomly choose to select $[E]$ to refine the library. Consequently, all factors of $[E]$ including $[E]$, $[S][E]$, $[P][E]$, $[ES][E]$, and $[E]^{2}$ are removed from the candidate library $\Theta_{0}$ (see how $\Theta_{1}$ is defined for $k=2$ in Figure 1b). This process of algebraic equation discovery and library refinement are iteratively continued until all algebraic relationships are discovered. Note that the algebraic equations found will not be unique because any linear combination of algebraic equations can lead to another algebraic relationship, but our approach tries to find a minimal set of equivalent algebraic relationships that can form any algebraic relationship using linear combinations.
@@
For many applications, such as power-grid networks, the number of algebraic constraints in a DAE system is known beforehand from physical principles such as the conservation of energy at each node. However, for other applications, such as CRNs, the number of algebraic constraints is not known in advance. When the number is unknown, an SVD analysis of the candidate library matrix $\Theta_{k-1}$ can be used to track the number of algebraic relationships remaining in the library (Figure 2) at iteration $k$. The right singular vectors (principal components) with zero singular values correspond to linear combinations in the library with zero variance (i.e., perfect multicollinearity), so the nullity of the covariance matrix estimates the number of algebraic relationships. The variance-explained ratio calculated for each singular value captures the relative importance of each principal direction in the library more clearly than the raw singular values. The ratio is the variance explained by each principal direction or singular value squared normalized by the total variance of the library or, equivalently, the sum of all singular values. When noise levels are minimal, there is a clear separation between the singular vectors lying in the nullspace and the dominant principal components. As the library is refined based on valid algebraic relationships, the zero singular values disappear while the number of dominant principal components remains constant until all zero singular values are removed (Figure 2a).
@@
Even at high noise, where the separation becomes less clear, the SVD spectrum can indicate whether the refinement is based on a true algebraic relationship in the library. For instance, at 15% noise in Figure 2b, refining the library using the term $[E]$ (which is part of an algebraic relationship) reduces the number of zero singular values while keeping the number of dominant principal directions—those with relatively large variance explained— unchanged. In contrast, refining it using $[P]$ (which is not part of any algebraic relationship) causes a loss in the number of dominant principal directions. Through a correct refinement, the candidate library can become significantly smaller and better-conditioned, aiding the discovery of dynamic equations. The effect of a series of correct library refinements in removing only low-variance-explained-ratio values is more pronounced for libraries with higher-order terms (e.g., monomials up to degree 4, see Figure 2c). If the number of algebraic relationships $K$ in the DAE system is not known beforehand, it can be treated either as a hyperparameter or determined via a Pareto front analysis of the library’s condition number improvements, which can serve as a stopping criterion for the algebraic discovery and refinement process ie, $\text{Cond}(\Theta_{K+1})\approx\text{Cond}(\Theta_{K})<\text{Cond}(\Theta_{K-1})<\cdots<\text{Cond}(\Theta_{0})$. By default, we use lack of improvement in condition number to determine $K$. For example, the improvement in conditioning stagnates after the first two refinements for the CRN example, and so the Pareto front suggests stopping library refinement after the second equation is removed (Figure 2d).
@@
Figure 2: SVD analysis of the candidate library $\Theta$: (a) library with monomials up to degree 2 and no noise, (b) library with monomials up to degree 2 and 15% noise, (c) library with monomials up to degree 4 and 15% noise, (d) tracking condition number after each refinement for library with monomials up to degree 4 and 15% noise. SVD analysis helps track redundancies in the candidate library space and determines when to stop iterations in the algebraic finder step.
In the DaeFinder package, the AlgModelFinder class provides many features to automate the steps in the algebraic finding step. This class supports various metrics for assigning scores to different algebraic relationships discovered as part of (13), including $\mathcal{R}^{2}$ values, mean squared error on test sets, etc. It also has functions to reduce the algebraic relationships and refine the candidate library based on SODAs.

### v2

@@ 2.3 Benefits and challenges of model discovery in the DAE framework
In addition to the high data quality requirements, a fundamental issue is that implicit-SINDy and SINDy-PI methods do not identify the algebraic variables and associated equations within a DAE system. The algebraic constraints in the DAEs typically carry physical significance related to physical constraints or quasi-steady-state approximations, and this intuitive meaning is lost in the equivalent ODE formulation. Creating an explicit system of ODEs in the form (8) can be complex, computationally intensive, or even impossible for higher-index DAE systems or when the algebraic constraints are not differentiable in time [Petzold [56]]. Furthermore, small parameter changes in the original DAE system often lead to distinct reduced ODE forms. Solving or discovering DAEs directly preserves the underlying structure of the problem. An additional benefit of identifying the algebraic relationships is that the library can be iteratively simplified by removing an algebraic variable for every algebraic equation discovered. Removal of terms with algebraic variables reduce multicollinearity among the candidate library terms due to the explicit algebraic constraints, improving the conditioning of the library. Consequently, the discovery of DAEs provides a more interpretable and computationally beneficial framing.
Discovering a DAE system in its semi-explicit form (2)–(3) also offers flexibility for numerical analysis and simulation. For example, in CRN applications, simulations are often carried out using ODE solvers designed for stiff systems; however, in applications involving high-index DAEs, solving a reduced ODE system can lead to numerical instability [Petzold [56], Kunkel [36], Brenan et al. [14]]. In these cases, single-step and multi-step methods for solving DAEs in their non-reduced forms are preferred [Petzold [56], Kunkel [36], Brenan et al. [14], Ascher and Petzold [4], Rabier and Rheinboldt [59]]. By first discovering the DAE system in the semi-explicit form, the user can choose either to convert it to a reduced ODE or to keep the original semi-explicit form, depending on which numerical solvers are most appropriate. In comparison, converting a reduced ODE system back into a DAE system with explicit constraints is often not feasible.
To address all the issues discussed so far, including the interpretability of physical constraints, the numerical instabilities caused by perfect multicollinearity in the candidate library, and the adversial effects of noise present in implicit ODE discovery, we propose the method SODAs. By separating the discovery of differential and algebraic equations, we directly identify the algebraic variables and equations and eliminate the impact of noisy derivatives on the discovery of algebraic equations. We can then utilize the knowledge gained from the discovered algebraic equations to better condition the library before attempting to discover differential equations.
@@ 3 Methods: Sparse Optimization for Differential-Algebraic Systems
Figure 1: SODAs algorithm explained using a chemical reaction network example. (a) Collect time series and construct candidate library matrix. (b) Algebraic Relation Finder: iteratively finds algebraic relationships and refines the candidate library based on the discovered relationship. In each iteration $k$, the best relationship is selected based on its score $S^{k}$, and a complexity score is assigned to each term within this relationship. This complexity score is used to select a term to refine the candidate library for the next iteration: the red X-s indicate the removal of features from the candidate library that are multiples of the selected term. (c) Dynamic Finder: the refined candidate library from the algebraic finder step is used to find the system of ODEs. (d) Assembled DAEs: The algebraic and differential equations from the algebraic and dynamic finder steps respectively, are assembled to form the final set of DAEs (left), which is equivalent to the true set of DAEs (right).
In this section, we introduce our algorithm SODAs (Figure 1, Algorithm 1), and demonstrate how it can be used to discover DAEs in their explicit form (2)–(3). Details on the inputs, parameters, and defaults used in Algorithm 1 can be found in Supplementary-section S1. The core idea is to separate the discovery of the algebraic and dynamic state equations. DaeFinder is a Python implementation of SODAs that allows users to easily create popular candidate libraries relevant to many applications and efficiently implement various components of the SODAs algorithm. Details about this package and defaults used in the algorithm can be found in Supplementary-section S2. SODAs iteratively identifies algebraic relationships, removes them from the library, and reduces multicollinearity within the library. After this process, SINDy is applied to discover the dynamic equations from the remaining library terms. We illustrate the application of SODAs using an example of a simple CRN (4)–(7) (Figure 1a).
@@
Let $\mathbf{X}=[\mathbf{X}_{1}\;\mathbf{X}_{2}\;\ldots\;\mathbf{X}_{d}]$ denote the $N\times d$ time-series data matrix, where each column $\mathbf{X}_{i}$ contains possibly noisy measurements of the state $x_{i}$ at time points $\{t_{1}=0,\ldots,t_{N}=T\}$, such that $\mathbf{X}_{ij}\approx x_{j}(t_{i})$. Note that the measurements $\textbf{X}_{ij}$ may also be obtained through transformations of the original state variables. However, for simplicity, throughout this paper, we assume that all states are directly observable with some measurement noise. SODAs is typically applied to DAE systems where the terms appearing in equations (2)–(3) can be written as linear combination of a set of predefined library terms which comes from prior knowledge. We start by defining a candidate library $\Theta=\{\theta_{1}(\mathbf{x}),\ldots,\theta_{J}(\mathbf{x})\}$, which is a set of functions of the state variables. Based on the time-series data $\mathbf{X}$, we next form the initial candidate library matrix denoted by $\Theta_{0}=[\theta_{1}(\mathbf{X})\ \theta_{2}(\mathbf{X})\ \ldots\ \theta_{J}(\mathbf{X})]$, which is an $N\times J$ matrix where the $j^{\text{th}}$ column represents the evaluation of the candidate function $\theta_{j}$ at time points $\{t_{1}=0,\ldots,t_{N}=T\}$. In our applications, a polynomial library is used for the CRN and the nonlinear pendulum, while a combination of polynomial and trigonometric functions is used for the power grid. After forming the candidate library, SODAs consists of two major steps: the algebraic finder step (Figure 1b) and the dynamical system finder step (Figure 1c).
@@
For many systems, it is unclear a priori which states are differential variables and which are algebraic, especially if separation of timescales is expected and a subset of variables are in quasi-steady state. Our approach does not necessarily require pre-existing knowledge about which states are in the algebraic relations. Notwithstanding, prior knowledge on state variables that take part in algebraic equations can be incorporated into the pipeline. For unknown algebraic equations, the algebraic finder step involves iteratively identifying algebraic relationships and reducing multicollinearity in the library by removing these relationships. Once all algebraic equations are found, the remaining dynamical equations are identified from the refined library using an existing method like SINDy [Brunton et al. [17], Wentz and Doostan [80], Messenger and Bortz [46]] or IDENT [Kang et al. [34], He et al. [29]].
@@
$L_{l}(\mathbf{p}^{l})$ $=\sum_{r=1}^{N}\Big(\theta_{l}(\mathbf{x}(t_{r}))-\sum_{\begin{subarray}{c}j\in\mathcal{I}_{k}\\ j\neq l\end{subarray}}p_{j}^{l}\theta_{j}(\mathbf{x}(t_{r}))\Big)^{2}+\alpha\|\mathbf{p}^{l}\|_{0},\quad l\in\mathcal{I}_{k-1}.$ (13)
@@
where $\Xi=\left[\mathbf{p}^{1}\ \mathbf{p}^{2}\ \ldots\ \mathbf{p}^{J_{k}}\right]$ with constraints $\Xi_{jj}=0$ and $\|\Xi\|_{0}$ is minimized (Figure 1b, left). Note that, in practice, the matrix form is not used, but instead each $\mathbf{p}^{l}$ is optimized in parallel by minimizing the loss function $L_{l}(\mathbf{p}^{l})$. A similar parallelized approach is introduced in [Kaheman et al. [33]] for semi-combinatorial discovery of implicit equations of the form (11), whereas in SODAs this parallel fitting approach is used for finding algebraic relationships. Notably, derivatives are not involved in the algebraic finder step and do not need to be estimated from the noisy data. There are multiple established approaches to enforce the sparsity of the parameter matrix $\Xi$. The hyperparameters used in this step include the thresholding parameter $\alpha_{\text{th}}$ [Brunton et al. [17]], which determines the coefficient size for removing terms, as well as the regularization penalty weight $\alpha$ [Tibshirani [75]], both of which are encoded in $\alpha^{a}$. In practice, these hyperparameters should be chosen using a Pareto-front approach, as discussed in [Brunton et al. [17]].
@@
The next step of the algorithm is to select the best-fitting algebraic relationship across all the candidate fits in (13) (Figure 1b, right). To identify the most likely relationship, we assign a score $S^{k}_{l}$ to each candidate fit in (13). In our examples, we have used the coefficient of determination $\mathcal{R}^{2}$ [James et al. [31]], which for a given feature $\theta_{l}$ with mean $\bar{\theta_{l}}$ is given by $\mathcal{R}^{2}=1-\frac{\sum_{r=1}^{N}\Big(\theta_{l}(\mathbf{x}(t_{r}))-\sum_{\begin{subarray}{c}j\in\mathcal{I}_{k}\\ j\neq l\end{subarray}}p_{j}^{l}\theta_{j}(\mathbf{x}(t_{r}))\Big)^{2}}{\sum_{r=1}^{N}\Big(\theta_{l}(\mathbf{x}(t_{r}))-\bar{\theta_{l}}\Big)^{2}}$, and ranges from 0 to 1, where larger values indicate a better fit. Other criteria, such as AIC and BIC (see [Mangan et al. [41], Chakrabarti and Ghosh [20]]), that balance training error with model simplicity can also be used to define $S^{k}_{l}$. The fit with the highest score is then identified as the most prominent algebraic relationship in the library $\Theta_{k-1}$ and is further used to refine $\Theta_{k-1}$. In our experience, conservation laws that involve first-order terms are initially identified as the best relationships during this process.
@@
After identifying the best algebraic relationship based on the score $S^{k}_{l}$ for the $k^{\text{th}}$ iteration, any common factors in that relationship are factored out to obtain a reduced equation $g_{k}(\mathbf{x})=0$. To address the multicollinearity introduced by this algebraic relationship, we remove one of the terms in $g_{k}(\mathbf{x})$. To decide which term to remove, each term $\theta^{(k)}$ in $g_{k}(\mathbf{x})$ is assigned a complexity score (Figure 1b, right). In polynomial libraries, this score is set to be the degree of the term by default, i.e., higher-degree monomials should have a higher complexity score. The term with the highest complexity score, $\theta^{(k)}_{\text{high }}$ (circled in Figure 1b, right), is then removed from the library. The rationale behind this refinement is that if $\theta^{(k)}_{\text{high }}$ is part of the discovered algebraic relationship $g_{k}(\mathbf{x})=0$, then $\theta^{(k)}_{\text{high }}$ can be expressed as a linear combination of the other terms in $g_{k}(\mathbf{x})$, so there is no loss of information by removing a term from the library. By keeping lower rather than higher complexity terms in the library, we reduce noise amplification [Messenger and Bortz [45]]. If multiple terms share the same complexity score, then either expert knowledge or quality of measurement of states (through statistical measures like variance) can also be used to choose one term over the other, or a term can be selected randomly. Additionally, for any $\theta\in\Theta_{k-1}$, the equation $\theta\,g_{k}(\mathbf{x})=0$ also represents a potential algebraic relationship that can lead to perfect multicollinearity. To remove all degenerate relationships of the form $\theta\,g_{k}(\mathbf{x})=0$, all multiples of $\theta^{(k)}_{\text{high }}$are also removed from $\Theta_{k-1}$. Let $\Gamma_{k}$ be a subset of the candidate library matrix $\Theta_{k-1}$ that has $\theta^{(k)}_{\text{high }}$ as a factor; then $\Theta_{k}=\Theta_{k-1}\setminus\Gamma_{k}$, where all elements of $\Gamma_{k}$ are removed from $\Theta_{k-1}$ as part of the refinement process. It follows that the libraries at each refinement step form a decreasing chain, $\Theta_{k}\subset\Theta_{k-1}\subset\cdots\subset\Theta_{0}.$
Mitigating multicollinearity due to algebraic relationships
@@
As an example, in the CRNs (4)–(7), the algebraic relationship $E_{1}]+[AE_{1}]-E_{1}^{tot}=0$ can cause the following perfect multicollinearity in the candidate library matrix $\Theta$ that has up to degree 2 monomials:
$[E_{1}]+[AE_{1}]-E_{1}^{tot}=0,$ (15)
$[E_{1}]\left([E_{1}]+[AE_{1}]-E_{1}^{tot}\right)=0,$ (16)
$[A]\left([E_{1}]+[AE_{1}]-E_{1}^{tot}\right)=0,$ (17)
$[B]\left([E_{1}]+[AE_{1}]-E_{1}^{tot}\right)=0,$ (18)
$[AE_{1}]\left([E_{1}]+[AE_{1}]-E_{1}^{tot}\right)=0.$ (19)
@@
For higher-order candidate libraries, there will be more relationships. All such degenerate algebraic relationships make up the nullspace of the library matrix $\Theta$. If any one of the degenerate equations (16)–(19) is identified as the relationship with the best $S_{l}^{k}$, it is first reduced to the form (15). Both terms in this reduced relationship (i.e., $E_{1}]$ and $AE_{1}]$) have the same complexity score of 1 based on the degree of the monomial (Figure 1b, top-right). In the absence of any other information, we randomly choose to select $E_{1}]$ to refine the library. Consequently, all factors of $E_{1}]$ including $E_{1}]$, $A][E_{1}]$, $B][E_{1}]$, $AE_{1}][E_{1}]$, and $E_{1}]^{2}$ are removed from the candidate library $\Theta_{0}$ (see how $\Theta_{1}$ is defined for $k=2$ in Figure 1b). This process of algebraic equation discovery and library refinement are iteratively continued until all algebraic relationships are discovered. Note that the algebraic equations found will not be unique because any linear combination of algebraic equations can lead to another algebraic relationship, but our approach tries to find a minimal set of equivalent algebraic relationships that can form any algebraic relationship using linear combinations.
@@
For many applications, such as power-grid networks, the number of algebraic constraints in a DAE system is known beforehand from physical principles such as the conservation of energy at each node. However, for other applications, such as CRNs, the number of algebraic constraints is not known in advance. When the number is unknown, an SVD analysis of the candidate library matrix $\Theta_{k-1}$ can be used to track the number of algebraic relationships remaining in the library (Figure 2) at iteration $k$. Consider the SVD decomposition
$\Theta_{k-1}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top},$ (20)
where $\boldsymbol{\Sigma}=\mathrm{diag}(\sigma_{1},\ldots,\sigma_{r})$ contains the singular values $\sigma_{1}\geq\cdots\geq\sigma_{r}\geq 0$, with $r=\mathrm{rank}(\Theta_{k-1})$. The right singular vectors associated with zero singular values ($\sigma_{i}\approx 0$) span the nullspace of $\Theta_{k-1}$ and correspond to exact linear dependencies in the library. Consequently, the nullity $\mathrm{nullity}(\Theta_{k-1})=J-r$ provides an estimate of the number of linear algebraic constraints that exist in the library. To quantify the relative contribution of each principal direction, we calculate the variance-explained ratio associated with the $i$-th singular value using
$\eta_{i}=\frac{\sigma_{i}^{2}}{\sum_{j=1}^{r}\sigma_{j}^{2}}.$ (21)
When noise levels are low, the spectrum $\{\eta_{i}\}$ shows a distinct separation between the dominant principal components and directions in the null space. As valid algebraic relationships are iteratively removed from the library, the zero singular values disappear, while the number of dominant components stays the same until all algebraic constraints have been eliminated. (Figure 2a).
@@
Even at high noise, where the separation becomes less clear, the SVD spectrum can indicate whether the refinement is based on a true algebraic relationship in the library. For instance, at 15% noise (see (22) to see how noise is added to time series) in Figure 2b, refining the library using the term $[E]$ (which is part of an algebraic relationship) reduces the number of zero singular values while keeping the number of dominant principal directions, those with relatively large variance explained, unchanged. In contrast, refining it using $[P]$ (which is not part of any algebraic relationship) causes a loss in the number of dominant principal directions. Through a correct refinement, the candidate library can become significantly smaller and better-conditioned, aiding the discovery of dynamic equations. The effect of a series of correct library refinements in removing only low-variance-explained-ratio values is more pronounced for libraries with higher-order terms (e.g., monomials up to degree 4, see Figure 2c). SODAs can be used with or without prior knowledge of the number of algebraic relationships, $K$. If $K$ in the DAE system is not known beforehand, it can determined via a Pareto front analysis of the library’s condition number improvements. We use the lack of improvement in the condition number within some tolerance as a stopping criterion for the algebraic discovery and refinement process ie, $\text{Cond}(\Theta_{K+1})\approx\text{Cond}(\Theta_{K})<\text{Cond}(\Theta_{K-1})<\cdots<\text{Cond}(\Theta_{0})$. For example, the improvement in conditioning stagnates after the first two refinements for the CRN example, and so the Pareto front suggests stopping library refinement after the second equation is removed (Figure 2d). For very noisy data it may be necessary to treat K as a hyperparameter. If some algebraic relationships are known ahead of time, then SODAs ensures at least that many relationships are recovered.
@@
Figure 2: SVD analysis of the candidate library $\Theta$: The Y-axes are presented on a natural logarithm scale. (a) library with monomials up to degree 2 and no noise, (b) library with monomials up to degree 2 and 15% noise, (c) library with monomials up to degree 4 and 15% noise, (d) tracking condition number after each refinement for library with monomials up to degree 4 and 15% noise. SVD analysis helps track redundancies in the candidate library space and determines when to stop iterations in the algebraic finder step.